### v1

@@ 2.2.2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q$ and a group of $K$-$V$ pairs to a new output vector. The output of the attention is a weighted sum of $V$ vectors, where the weight for each $V$ is computed by taking the dot product of the $Q$ with the corresponding $K$ vaswani2017attention . In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation sam to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of sam can be described as follows:
@@
Each $Q$, $K$, and $V$ is partitioned into $p$ parts (called heads), denoted by $Q^{i}$,$K^{i}$, and $V^{i}$, $i=1,2,\ldots,p$, and the dimension of each part (i.e., the number of columns) is $k^{\prime}=k/p$. We use $Q^{q}_{i}$, $K^{q}_{i}$, and $V^{q}_{i}$ to denote the feature of node $i$ in $Q^{q}$, $K^{q}$, and $V^{q}$ (i.e., the row vector corresponding to $i$), respectively. Let $X^{q}$ be the $q$-th head of $X\in\{Q,K,V\}$. Then, the weight of aggregating neighbor information is calculated based on $Q^{q}$ and $K^{q}$. Specifically, for a node $i$ (drug or disease), the weight $\beta^{r}_{ij}$ of $i$ aggregating its neighbor $j$ is defined as
@@
$H_{\texttt{sam}}^{i}=\texttt{concat}(\sum_{j\in N(i)}\beta^{1}_{ij}V^{1}_{j},\sum_{j\in N(i)}\beta^{2}_{ij}V^{2}_{j},\ldots,\sum_{j\in N(i)}\beta^{p}_{ij}V^{p}_{j})$ (5)
Moreover, let $H_{\texttt{sam}}$ represent the feature matrix obtained by SAM, $H_{r}\in\mathbb{R}^{n\times k}$ be the drug feature obtained by SAM, and $H_{d}\in\mathbb{R}^{m\times k}$ is the disease feature obtained by SAM. Then, we have
$H_{r}=\begin{bmatrix}H^{r_{1}}_{\texttt{sam}}\\ H^{r_{2}}_{\texttt{sam}}\\ \dots\\ H^{r_{m}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\texttt{sam}}\\ H^{d_{2}}_{\texttt{sam}}\\ \dots\\ H^{d_{n}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{\texttt{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
A fully connected layer is applied after the SAM function to extract further features of drugs and diseases. We use the notation samf to denote the whole process, i.e., samf is the combination of sam with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU maas2013leakyrelu activation function. Then,
$\texttt{samf}(E,H^{in})$ $=\texttt{FC}(\texttt{sam}(E,H^{in}))$ (7)
@@ 2.2.3 Encoder
Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$. In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\texttt{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE. Then,
@@
$=(\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
For the InterDDFE module, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer. Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\texttt{samf}^{\ell}_{g}$ represents the aggregation function. Then,
@@
${=(\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
It should be noted that compared with the IntraDDFE module, the InterDDFE module changes the features: the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$ and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
@@ 2.2.4 Decoder
${s_{r}}$ $=\texttt{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
${s_{d}}$ $=\texttt{sigmod}(H^{d}_{a}\times(H^{r}_{s})^{T})$ (16)
@@ 2.3 Optimization
Predicting the drug-disease association can be viewed as a binary classification task. The objective is to determine whether a given drug is associated with a specific disease. Given the serious imbalance between the number of positive samples (drug-disease associations) and negative samples (drug-disease non-associations), we employ binary cross-entropy as the loss function. We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively. We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($i$)-disease($j$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
$loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{i,j})+\sum_{(i,j)\in y^{-}}\log(1-s_{i,j}))$ (18)
We optimize the DFDRNN model using the Adam kingma2014adam optimizer and initialize the parameters’ weights with Xavier initialization glorot2010understanding . We randomly drop some edges and apply regular dropout srivastava2014dropout to improve the model’s generalization ability.
@@ 3.1 Baseline methods
$\bullet$ LAGCN lagcn : This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN cai2021DRHGCN : This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF meng2022DRWBNCF : This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA zhang2024NCH-DDA : This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT huang2024HDGAT : This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN tang2023DRGBCN : This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
@@ 3.3 Evaluation metrics
Table 2: The statistics of the four datasets used in this study
@@ 3.4 Performance of DFDRNN in the cross-validation
[…]
@@
[…]
{tablenotes}
The best results in each row are in bold faces and the second best results are underlined.
@@ 3.5 Effect of the number of nearest neighbors
Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$. However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention aggregation capability. Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across four datasets. For detailed numerical values, please refer to Supplementary Tables S1 and S2.
@@ 3.6.2 Decoder
${s^{non}_{r}}$ $=\texttt{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
${s^{non}_{d}}$ $=\texttt{sigmod}(H^{d}_{s}\times(H^{r}_{s})^{T})$ (23)
@@ 3.7 Identification of candidates for new diseases
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic singh2020drug . Finding suitable drugs for emerging diseases is a complex and crucial task. To evaluate DFDRNN’s predictive ability in identifying potential candidate drugs for new diseases, we conducted leave-one-out cross-validation (LOOCV) on the Fdataset dataset. Specifically, for each disease $d_{x}$, all drug-disease associations related to $d_{x}$ were excluded from the analysis and treated as test samples. The remaining drug-disease associations were used as training samples, relying solely on the similarity between disease $d_{x}$ and other diseases to predict its candidate drugs.
@@
In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated. The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCHDDA by 0.99%. Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
@@ 3.8 Case study
To assess the real-world performance of DFDRNN, we conducted case studies on two diseases (Alzheimer’s disease (AD) and Parkinson’s disease (PD)) from the Fdataset. We utilized all drug-disease associations in the Fdataset as the training set to train DFDRNN and predict the association probabilities for all unknown drug-disease pairs. Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD. To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB knox2024drugbank , PubChem kim2016pubchem , CTD davis2023CTD , DrugCentral avram2021drugcentral , and ClinicalTrials.
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table 5 displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment erkulwater1989amantadine . Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients lemke1995effect . DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
@@
[…]
@@ 10 DB01262 Decitabine CTD
[…]

### v2

@@ 2.2.2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector. The output of the attention is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ (Vaswani et al., 2017). In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of $\mathbf{sam}$ can be described as follows:
@@
Each $Q$, $K$, and $V$ is partitioned into $p$ parts (called heads), denoted by $Q^{i}$,$K^{i}$, and $V^{i}$, $i=1,2,\ldots,p$, and the dimension of each part (i.e., the number of columns) is $k^{\prime}=k/p$. We use $Q^{q}_{i}$, $K^{q}_{i}$, and $V^{q}_{i}$ to denote the feature of node $i$ in $Q^{q}$, $K^{q}$, and $V^{q}$ (i.e., the row vector corresponding to $i$), respectively. Let $X^{q}$ be the $q$-th head of $X\in\{Q,K,V\}$. Then, the weight of aggregating neighbor information is calculated based on $Q^{q}$ and $K^{q}$. Specifically, for a node $i$ (drug or disease), the weight $\beta^{q}_{ij}$ of $i$ aggregating its neighbor $j$ is defined as
@@
$H_{\mathbf{sam}}^{i}=\mathbf{concat}(\sum_{j\in N(i)}\beta^{1}_{ij}V^{1}_{j},\sum_{j\in N(i)}\beta^{2}_{ij}V^{2}_{j},\ldots,\sum_{j\in N(i)}\beta^{p}_{ij}V^{p}_{j})$ (5)
Moreover, let $H_{\mathbf{sam}}$ represent the feature matrix obtained by SAM, $H_{r}\in\mathbb{R}^{n\times k}$ be the drug feature obtained by SAM, and $H_{d}\in\mathbb{R}^{m\times k}$ is the disease feature obtained by SAM. Then, we have
$H_{r}=\begin{bmatrix}H^{r_{1}}_{\mathbf{sam}}\\ H^{r_{2}}_{\mathbf{sam}}\\ \dots\\ H^{r_{n}}_{\mathbf{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\mathbf{sam}}\\ H^{d_{2}}_{\mathbf{sam}}\\ \dots\\ H^{d_{m}}_{\mathbf{sam}}\end{bmatrix},\hskip 8.5359ptH_{\mathbf{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
A fully connected layer is applied after the SAM function to extract further features of drugs and diseases. We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU (Maas et al., 2013) activation function. Then,
$\mathbf{samf}(E,H^{in})$ $=\mathbf{FC}(\mathbf{sam}(E,H^{in}))$ (7)
@@ 2.2.3 Encoder
Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$. In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\mathbf{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE. Then,
@@
$=(\mathbf{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\mathbf{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
Since IntraDDFE extracts features within a single domain, the dual-feature does not undergo dynamic changes. Thus, the input similarity features $H^{\ell}_{s}$ produce the output similarity features $\hat{H}^{\ell+1}_{s}$, and the input association features $H^{\ell}_{a}$ produce the output association features $\hat{H}^{\ell+1}_{a}$. Next, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer. Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\mathbf{samf}^{\ell}_{g}$ represents the aggregation function. Then,
@@
${=(\mathbf{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\mathbf{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
Since the InterDDFE module extracts features across domains, dual-feature undergo dynamic changes. Specifically, the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$, and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
@@ 2.2.4 Decoder
${s_{r}}$ $=\mathbf{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
${s_{d}}$ $=\mathbf{sigmod}(H^{d}_{a}\times(H^{r}_{s})^{T})$ (16)
@@ 2.3 Optimization
Predicting the drug-disease association can be viewed as a binary classification task. The objective is to determine whether a given drug is associated with a specific disease. Given the serious imbalance between the number of positive samples (drug-disease associations) and negative samples (drug-disease non-associations), we employ binary cross-entropy as the loss function. We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively. We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($r_{i}$)-disease($d_{j}$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
$loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{ij})+\sum_{(i,j)\in y^{-}}\log(1-s_{ij}))$ (18)
We optimize the DFDRNN model using the Adam (Kingma and Ba, 2014) optimizer and initialize the parameters’ weights with Xavier initialization (Glorot and Bengio, 2010). We randomly drop some edges and apply regular dropout (Srivastava et al., 2014) to improve the model’s generalization ability.
@@ 3.1 Baseline methods
$\bullet$ LAGCN (Yu et al., 2021): This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN (Cai et al., 2021): This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF (Meng et al., 2022): This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA (Zhang et al., 2024): This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT (Huang et al., 2024): This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN (Tang et al., 2023): This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
@@ 3.3 Evaluation metrics
Table 2: The statistics of the four sets used in this study
@@ 3.4 Performance of DFDRNN in the cross-validation
Metrics Model Fdataset Cdataset Ldataset LRSSL Avg.
@@
  • The best results in each column are in bold faces and the second best results are underlined.
@@ 3.5 Effect of the number of nearest neighbors
Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$. However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention mechanism’s aggregation capability. Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across the four datasets. For detailed numerical values, please refer to Supplementary Tables S1 and S2.
@@ 3.6.2 Decoder
${s^{non}_{r}}$ $=\mathbf{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
${s^{non}_{d}}$ $=\mathbf{sigmod}(H^{d}_{s}\times(H^{r}_{s})^{T})$ (23)
@@ 3.7 Identification of candidates for new diseases
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic (Singh et al., 2020). Finding suitable drugs for emerging diseases is a complex and crucial task. To evaluate DFDRNN’s predictive ability in identifying potential candidate drugs for new diseases, we conducted leave-one-out cross-validation (LOOCV) on the Fdataset dataset. Specifically, for each disease $d_{x}$, all drug-disease associations related to $d_{x}$ were excluded from the analysis and treated as test samples. The remaining drug-disease associations were used as training samples, relying solely on the similarity between disease $d_{x}$ and other diseases to predict its candidate drugs.
@@
In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated. The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCH-DDA by 0.99%. Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
@@ 3.8 Case study
To assess the real-world performance of DFDRNN, we conducted case studies on two diseases (Alzheimer’s disease (AD) and Parkinson’s disease (PD)) from the Fdataset. We utilized all drug-disease associations in the Fdataset as the training set to train DFDRNN and predict the association probabilities for all unknown drug-disease pairs. Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD. To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB (Knox et al., 2024), PubChem (Kim et al., 2016), CTD (Davis et al., 2023), DrugCentral (Avram et al., 2021), and ClinicalTrials.
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table 5 displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment (Erkulwater and Pillai, 1989). Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients (Lemke, 1995). DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
@@
Rank Drug ID Drug Name Evidence
1 DB00915 Amantadine DB/CTD/DrugCentral/ PubChem/ClinicalTrials
@@ 10 DB01262 Decitabine CTD
Rank Drug ID Drug Name Evidence
1 DB00502 Haloperidol DB/CTD/PubChem/ ClinicalTrials